OpenStudy (anonymous):

what is the Slope- intercept form for points (4,4), (7,4)

OpenStudy (anonymous):

@DanJS

OpenStudy (solomonzelman):

you can tell it is a horizontal line, correct?

OpenStudy (danjs):

can you plot those 2 points on a graph?

OpenStudy (danjs):

Slope = (4-4) / (7-4) = 0 / 3 = 0 a line with zero slope, will be horizontal

OpenStudy (anonymous):



OpenStudy (danjs):

right on, so the slope is zero, any value of x will be 4

OpenStudy (danjs):

in short y = 4 is enough to describe the line

OpenStudy (anonymous):

yeh so then what is the slope- intercept form tho

OpenStudy (danjs):

if you choose a point say (4,4) and the slope = 0 y - y1 = m*(x-x1) y - 4 = 0(x - 4) y - 4 = 0 y = 4

OpenStudy (danjs):

for all values of x, y=4

OpenStudy (anonymous):

ok so does that mean y=4 is the slope- intercept form?

OpenStudy (danjs):

well it is the same as saying y = 0x + 4